Ticket Number: HEA-NOC/20040112-7 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20040112 16:18 UTC Problem Start: 20040114 08:00 UTC
Ticket Closed: 20040115 10:45 UTC Problem End: 20040114 09:00 UTC
To improve our co-location service the connecting switch (Cisco 2924) by an
upgraded switch (Cisco Cat3750). This will require switchover of all
connected servers; linux.ie, HEA webserver, E-Grants system, EPA.
New switch allows for GE uplink to HEAnet Brooklawn House PoP as well as
allowing for Gigabit connectivity to our co-located customers. There will
be a 30 second outage per customer during the switchover.
20040112 16:09 BN Switch has been configured by Colin and is in place.
Switchover of connections due to take place between 08:00 and 09:00
Wednesday 14th January.
20040114 08:54 BN All changeovers complete, all systems up. Switch will
be monitored for 24 hours.
20040115 10:45 BN Switch has been stable for over 24 hours.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20040112-7